# On the EFT validity for Drell-Yan tails at the LHC

### v1

@@ Abstract
preprint: TTK-24-53preprint: P3H-24-094
@@ I Introduction
The main caveat of EFT analyses at colliders is that the experimental sensitivity on $\mathcal{C}/\Lambda^{2}$ is not always sufficient to consistently probe EFTs, i.e., with $E\ll\Lambda$, while having Wilson coefficients $\mathcal{C}$ within the perturbative regime Brivio:2022pyi . The applicability range of an EFT and the impact of $d=8$ operators can be estimated in various manners Contino:2016jqw ; Azatov:2016sqh ; Cohen:2021gdw ; Brivio:2022pyi . For instance, by defining the so-called maximal cutoff scale ($\Lambda_{\mathrm{max}}$) that can be consistently probed for a perturbative scenario. The value of $\Lambda_{\mathrm{max}}$ is defined by the requirement that the $2\to 2$ scattering amplitude does not exceed the $16\pi^{2}$ limit arising from perturbativity Farina:2016rws , giving a first indication of the limits of the EFT description. Another approach often used is to provide constraints as a function of a mass scale $M_{\text{cut}}$ in some kinematical variable relevant to the process considered, above which all data are discarded Contino:2016jqw ; Farina:2016rws . In this way, it is possible to extract weaker, but potentially more robust, EFT limits, even though the phase-space migration of events in the detector simulation can be a potential problem Falkowski:2016cxu ; Brivio:2022pyi .
@@ IV.1 Partonic description
We turn now our attention to the $U_{1}$ leptoquark model defined in Eq. (18). The ratio of the full $d_{i}\bar{d}_{j}\to\ell_{\alpha}^{-}\ell_{\alpha}^{+}$ cross section with respect to the SM one can be written in a similar way after integration over $\hat{t}\in(-\hat{s},0)$. Neglecting the leptoquark width, we can write Jaffredo:2021ymt ; Davidson:2013fxa ; Davidson:2014lsa ; Eboli:1987vb )
@@
Figure 1: The ratio of the differential cross section in the EFT approach ($\mathrm{d}\hat{\sigma}_{\mathrm{EFT}}/\mathrm{d}m_{\ell\ell}^{2}$) and the $Z^{\prime}\sim(\mathbf{1},\mathbf{1},0)$ model ($\mathrm{d}\hat{\sigma}_{Z^{\prime}}/\mathrm{d}m_{\ell\ell}^{2}$) are plotted as a function of $\smash{m_{\ell\ell}^{2}/m_{Z^{\prime}}^{2}}$, where $m_{\ell\ell}=\sqrt{\hat{s}}$ is the invariant mass of the dilepton system. The mediator mass is fixed to $m_{Z^{\prime}}=2~{}\mathrm{TeV}$ and the couplings to $\smash{g^{(l)}_{\ell\ell}\,g^{(q)}_{ii}=1}$ (upper row) and $\smash{g^{(l)}_{\ell\ell}\,g^{(q)}_{ii}=(0.3)^{2}}$ (lower row), for fixed lepton flavors, d and different quark flavors in the columns of this plot. The EFT cross section is computed at different orders in the EFT expansion, with contributions up to interference of $d=6$ terms with the SM (dashed green), squared $d=6$ terms (solid green), interference of $d=8$ terms with the SM (dashed purple), and squared $d=8$ terms (solid purple).
@@ IV.2 Cross sections and quark flavor dependence
In Fig. 1, we compare the EFT cross sections to the predictions of the full $Z^{\prime}$ model, as a function of the ratio $m_{\ell\ell}^{2}/m_{Z^{\prime}}^{2}$, for a fixed mass $m_{Z^{\prime}}=2~{}\mathrm{TeV}$ and fixed lepton flavor $\ell$, where $m_{\ell\ell}=\smash{\sqrt{\hat{s}}}$ is the dilepton invariant mass. Specifically, the ratio of the differential EFT cross section, computed up to different truncation orders, to that obtained in the full model is plotted. The dashed green line is determined only taking into account the interference of $d=6$ EFT operators with the SM amplitude, while the solid green, dashed purple, and solid purple lines sequentially add contributions from the square of $d=6$ terms, the interference of $d=8$ operators with the SM, and the square of the $d=8$ terms, respectively. Notice that the dashed lines correspond to truncating the EFT series on cross-section level at ${\mathcal{O}}_{\!\sigma}(\Lambda^{-2})$ and ${\mathcal{O}}_{\!\sigma}(\Lambda^{-4})$ for the green and purple lines, respectively. The solid lines, on the other hand, represent a truncation on amplitude level at ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-2})$ and ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-4})$ again for the green and purple lines, respectively. Here, we have introduced the notation ${\mathcal{O}}_{\!\sigma}(\Lambda^{-n})$ and ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ to indicate a truncation of the EFT series on the cross-section and amplitude level, respectively. A further discussion of the two truncation approaches is presented in Sec. IV.4. To obtain the EFT cross section, the matching conditions in Eqs. (16) and (17) are employed. The light gray dashed line indicates the cross section of the full model. The couplings are fixed to $\smash{g^{(\ell)}g^{(q)}=1}$ (upper row) and $\smash{g^{(\ell)}g^{(q)}=(0.3)^{2}}$ (lower row), with flavor-diagonal couplings to quarks and leptons. The flavor of the quarks to which the $Z^{\prime}$ is coupled is varied from the first (left column) to the second (center column) and third (right column) generation. A similar comparison is made in Fig. 2 for the $U_{1}$ leptoquark model, again considering two fixed values of the couplings $x_{L}^{q\ell}$ and different quark flavors.
@@ IV.3 Discussion
where ${\mathcal{A}}_{\mathrm{SM}},\,{\mathcal{A}}_{6},\,{\mathcal{A}}_{6\times 6}$, and ${\mathcal{A}}_{8}$ represent the amplitudes of the SM, the EFT with a single $d=6$ operator insertion, with a double $d=6$ operator insertion, and a single $d=8$ operator insertion, respectively. The EFT power counting is made explicit here by factoring out the appropriate powers of the EFT cutoff scale $\Lambda$. The ellipses indicate terms of order ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-6})$ in the EFT expansion of the amplitude. For low-energy measurements (e.g., in the flavor sector) considering only the linear $d=6$ interference terms of ${\mathcal{O}}_{\!\sigma}(\Lambda^{-2})$ suffices and the $(d=6)^{2}$ contributions of ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-2})$ can be neglected due to the large scale suppression. Investigating the EFT truncation for high-energy observables, such as the Drell-Yan tails, is more subtle. The shapes of distributions for high-$p_{T}$ observables are determined by the high-energy properties of the corresponding amplitudes, which in turn are governed by their analytical structure and the principle of unitarity Adams:2006sv , even though the latter is lost for the EFT at very high energies $E>\Lambda$ above the cutoff.  [^8] Notice, in particular, that the Wilson coefficients of $d=8$ operators of the considered class ${\mathcal{O}}_{\psi^{4}D^{2}}$ are also subject to positivity constraints Chala:2023xjy (see also Ref. Remmen:2019cyz ). Here, however, we consider energies where the unitarity is not yet lost. When truncating the EFT series on amplitude level ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ for a given $n\in\mathbb{N}$, these properties are retained below the cutoff $E<\Lambda$. However, when truncating on cross-section level ${\mathcal{O}}_{\!\sigma}(\Lambda^{-n})$ these properties are upset since this does not correspond to a consistent truncation of the amplitude. In the latter case, even negative values for cross sections are, in principle, possible.
@@
Furthermore, we notice that the EFT series is unphysical. This can be understood through the LSZ formula Lehmann:1954rq which permits the freedom to perform field redefinitions for a theory without changing its $S$ matrix, that is, physical observables. However, in the case of an EFT, these field redefinitions generally shift contributions between different orders in the EFT expansions, rendering the expansion unphysical. Therefore, it appears that requiring a consistent EFT truncation for physical observables is unwarranted. Instead, it is useful to adopt a top-down perspective to determine a consistent EFT truncation prescription for high-energy observables. While the EFT is useful for studying experimental data in a model-independent framework, it should ultimately always be matched onto concrete underlying UV models of interest. When this is done, the EFT is constructed (or in other words, its coefficients are adjusted) in such a way that the resulting EFT amplitudes are mimicking the corresponding UV amplitude to a given accuracy in the EFT power counting, see e.g. Ref. DeAngelis:2023bmd 999 This is the case for both off-shell and on-shell matching, where the corresponding EFT amplitudes are equated with the ones of the full UV model in order to determine the matching conditions. This naturally suggests that also in a bottom-up scenarios, the EFT series should be truncated on amplitude rather than cross-section level. Only the former guarantees that the EFT properly reproduces the behavior of UV models. We therefore suggest truncating the EFT series on amplitude level ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ for some $n\in\mathbb{N}$. As explicitly verified for Drell-Yan processes in Figs. 1 and 2 with $n=2$ and $4$ the ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ truncations (solid lines) provide better convergence toward the full model prediction than the cross-section truncation ${\mathcal{O}}_{\!\sigma}(\Lambda^{-n})$ (dashed lines). According to this prescription, the leading order ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-2})$ EFT contributions contain both interference of $d=6$ EFT operators with the SM and the $(d=6)^{2}$ contributions to the cross section, which can also be seen in Eq. (31). As discussed in Ref. Brivio:2022pyi this provides a well-defined, unambiguous, and gauge-invariant setup. At the first subleading order ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-4})$ interference of $d=8$ operators with the SM, $(d=8)^{2}$ contributions, as well as the interference of $(d=6)$ with $(d=8)$ operators are included. In addition to dimension-eight operators contributing to the amplitude, the corresponding contributions from double insertion of two $d=6$ operators into the same amplitude are taken into account at this order as well, cf. Eq. (31), although this is not relevant for the present discussion of Drell-Yan tails at tree level.
@@ VI Illustration: $b\to c\tau\nu$
where we have again assumed down alignment for the left-handed quark doublet, and kept couplings to the $\tau$ lepton, and the second- and third-generation quarks. Other relevant constraints on this scenario are posed by the modifications of $Z$ and $W$ couplings to leptons induced by the Renormalization Group (RG) evolution, which are sizable as they are proportional to the top Yukawa coupling $y_{t}$ in this case Feruglio:2016gvd ; Feruglio:2017rjo ; Cornella:2018tfd ; Allwicher:2022vbf ; Allwicher:2023aql . In our case, the biggest effect comes from $Z\nu\bar{\nu}$ vertex corrections, parametrized by Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga
@@
In summary, we find that EFTs can reliably describe Drell-Yan tails for $t$- and $u$-channel mediators with masses above $1.5$ TeV, including the models proposed to address the $R_{D^{(\ast)}}$ anomalies. We have shown that the EFT description at $\mathcal{O}_{\mathcal{A}}(\Lambda^{-2})$ is rather accurate for these masses and it can be systematically improved by including the $\mathcal{O}_{\mathcal{A}}(\Lambda^{-4})$ contributions which are already implemented in HighPT Allwicher:2022mcg .

### v2

@@ Abstract
preprint: TTK-24-53preprint: P3H-24-094preprint: DESY-24-217
@@ I Introduction
The main caveat of EFT analyses at colliders is that the experimental sensitivity on $\mathcal{C}/\Lambda^{2}$ is not always sufficient to consistently probe EFTs, i.e., with $E\ll\Lambda$, while having Wilson coefficients $\mathcal{C}$ within the perturbative regime Brivio:2022pyi . The applicability range of an EFT and the impact of $d=8$ operators can be estimated in various manners Contino:2016jqw ; Azatov:2016sqh ; Cohen:2021gdw ; Brivio:2022pyi ; Corbett:2024evt . For instance, by defining the so-called maximal cutoff scale ($\Lambda_{\mathrm{max}}$) that can be consistently probed for a perturbative scenario. The value of $\Lambda_{\mathrm{max}}$ is defined by the requirement that the $2\to 2$ scattering amplitude does not exceed the $16\pi^{2}$ limit arising from perturbativity Farina:2016rws , giving a first indication of the limits of the EFT description. Another approach often used is to provide constraints as a function of a mass scale $M_{\text{cut}}$ in some kinematical variable relevant to the process considered, above which all data are discarded Contino:2016jqw ; Farina:2016rws . In this way, it is possible to extract weaker, but potentially more robust, EFT limits, even though the phase-space migration of events in the detector simulation can be a potential problem Falkowski:2016cxu ; Brivio:2022pyi .
@@ IV.1 Partonic description
We turn now our attention to the $U_{1}$ leptoquark model defined in Eq. (18). The ratio of the full $d_{i}\bar{d}_{j}\to\ell_{\alpha}^{-}\ell_{\alpha}^{+}$ cross section with respect to the SM one can be written in a similar way after integration over $\hat{t}\in(-\hat{s},0)$. Neglecting the leptoquark width, we can write Jaffredo:2021ymt ; Davidson:2013fxa ; Davidson:2014lsa ; Eboli:1987vb
@@
Figure 1: The ratio of the differential cross section in the EFT approach ($\mathrm{d}\hat{\sigma}_{\mathrm{EFT}}/\mathrm{d}m_{\ell\ell}^{2}$) and the $Z^{\prime}\sim(\mathbf{1},\mathbf{1},0)$ model ($\mathrm{d}\hat{\sigma}_{Z^{\prime}}/\mathrm{d}m_{\ell\ell}^{2}$) are plotted as a function of $\smash{m_{\ell\ell}^{2}/m_{Z^{\prime}}^{2}}$, where $m_{\ell\ell}=\sqrt{\hat{s}}$ is the invariant mass of the dilepton system. The mediator mass is fixed to $m_{Z^{\prime}}=2~{}\mathrm{TeV}$ and the couplings to $\smash{g^{(l)}_{\ell\ell}\,g^{(q)}_{ii}=1}$ (upper row) and $\smash{g^{(l)}_{\ell\ell}\,g^{(q)}_{ii}=(0.3)^{2}}$ (lower row), for fixed lepton flavors and different quark flavors in the columns of this plot. The EFT cross section is computed at different orders in the EFT expansion, with contributions up to interference of $d=6$ terms with the SM (dashed green), squared $d=6$ terms (solid green), interference of $d=8$ terms with the SM (dashed purple), and squared $d=8$ terms (solid purple).
@@ IV.2 Cross sections and quark flavor dependence
In Fig. 1, we compare the EFT cross sections to the predictions of the full $Z^{\prime}$ model, as a function of the ratio $m_{\ell\ell}^{2}/m_{Z^{\prime}}^{2}$, for a fixed mass $m_{Z^{\prime}}=2~{}\mathrm{TeV}$ and fixed lepton flavor $\ell$, where $m_{\ell\ell}=\smash{\sqrt{\hat{s}}}$ is the dilepton invariant mass.  [^8] The full cross sections have been computed at tree level using the HighPT package, not relying on the expansions shown in Eqs. (26) and (30). Specifically, the ratio of the differential EFT cross section, computed up to different truncation orders, to that obtained in the full model is plotted. The dashed green line is determined only taking into account the interference of $d=6$ EFT operators with the SM amplitude, while the solid green, dashed purple, and solid purple lines sequentially add contributions from the square of $d=6$ terms, the interference of $d=8$ operators with the SM, and the square of the $d=8$ terms, respectively. Notice that the dashed lines correspond to truncating the EFT series on cross-section level at ${\mathcal{O}}_{\!\sigma}(\Lambda^{-2})$ and ${\mathcal{O}}_{\!\sigma}(\Lambda^{-4})$ for the green and purple lines, respectively. The solid lines, on the other hand, represent a truncation on amplitude level at ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-2})$ and ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-4})$ again for the green and purple lines, respectively. Here, we have introduced the notation ${\mathcal{O}}_{\!\sigma}(\Lambda^{-n})$ and ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ to indicate a truncation of the EFT series on the cross-section and amplitude level, respectively. A further discussion of the two truncation approaches is presented in Sec. IV.4. To obtain the EFT cross section, the matching conditions in Eqs. (16) and (17) are employed. The light gray dashed line indicates the cross section of the full model. The couplings are fixed to $\smash{g^{(\ell)}g^{(q)}=1}$ (upper row) and $\smash{g^{(\ell)}g^{(q)}=(0.3)^{2}}$ (lower row), with flavor-diagonal couplings to quarks and leptons. The flavor of the quarks to which the $Z^{\prime}$ is coupled is varied from the first (left column) to the second (center column) and third (right column) generation. A similar comparison is made in Fig. 2 for the $U_{1}$ leptoquark model, again considering two fixed values of the couplings $x_{L}^{q\ell}$ and different quark flavors.
@@ IV.3 Discussion
Note also that we neglect Renormalization Group (RG) evolution effects between the different energy scales probed by the Drell-Yan data. Heavy new physics scenarios are mostly sensitive to the high-$p_{T}$ tails of these distributions where RG effects are small. Moreover, the RG effects are similar between a UV theory and its corresponding low-energy EFT. Therefore, when the ratio of the cross sections is taken, we expect these effects to largely cancel out and not affect the study of the EFT validity. However, for a broader phenomenological analysis (which is beyond the scope of the present work), these effects should be considered Bartocci:2024fmm .
@@
where ${\mathcal{A}}_{\mathrm{SM}},\,{\mathcal{A}}_{6},\,{\mathcal{A}}_{6\times 6}$, and ${\mathcal{A}}_{8}$ represent the amplitudes of the SM, the EFT with a single $d=6$ operator insertion, with a double $d=6$ operator insertion, and a single $d=8$ operator insertion, respectively. The EFT power counting is made explicit here by factoring out the appropriate powers of the EFT cutoff scale $\Lambda$. The ellipsis indicate terms of order ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-6})$ in the EFT expansion of the amplitude. For low-energy measurements (e.g., in the flavor sector) considering only the linear $d=6$ interference terms of ${\mathcal{O}}_{\!\sigma}(\Lambda^{-2})$ suffices and the $(d=6)^{2}$ contributions of ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-2})$ can be neglected due to the large scale suppression. Investigating the EFT truncation for high-energy observables, such as the Drell-Yan tails, is more subtle. The shapes of distributions for high-$p_{T}$ observables are determined by the high-energy properties of the corresponding amplitudes, which in turn are governed by their analytical structure and the principle of unitarity Adams:2006sv , even though the latter is lost for the EFT at very high energies $E>\Lambda$ above the cutoff.  [^9] Notice, in particular, that the Wilson coefficients of $d=8$ operators of the considered class ${\mathcal{O}}_{\psi^{4}D^{2}}$ are also subject to positivity constraints Chala:2023xjy (see also Ref. Remmen:2019cyz ). Here, however, we consider energies where the unitarity is not yet lost. When truncating the EFT series on amplitude level ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ for a given $n\in\mathbb{N}$, these properties are retained below the cutoff $E<\Lambda$. However, when truncating on cross-section level ${\mathcal{O}}_{\!\sigma}(\Lambda^{-n})$ these properties are upset since this does not correspond to a consistent truncation of the amplitude. In the latter case, even negative values for cross sections are, in principle, possible.
@@
Furthermore, we notice that the EFT series is unphysical. This can be understood through the LSZ formula Lehmann:1954rq which permits the freedom to perform field redefinitions for a theory without changing its $S$ matrix, that is, physical observables. However, in the case of an EFT, these field redefinitions generally shift contributions between different orders in the EFT expansions, rendering the expansion unphysical. Therefore, it appears that requiring a consistent EFT truncation for physical observables is unwarranted. Instead, it is useful to adopt a top-down perspective to determine a consistent EFT truncation prescription for high-energy observables. While the EFT is useful for studying experimental data in a model-independent framework, it should ultimately always be matched onto concrete underlying UV models of interest. When this is done, the EFT is constructed (or in other words, its coefficients are adjusted) in such a way that the resulting EFT amplitudes are mimicking the corresponding UV amplitude to a given accuracy in the EFT power counting, see e.g. Ref. DeAngelis:2023bmd 101010 This is the case for both off-shell and on-shell matching, where the corresponding EFT amplitudes are equated with the ones of the full UV model in order to determine the matching conditions. This naturally suggests that also in a bottom-up scenarios, the EFT series should be truncated on amplitude rather than cross-section level, at least if the bottom-up analysis is intended to be linked to realistic UV scenarios at a later stage.  [^11] In a pure bottom-up setup, where there is no intention of linking to UV theories, this strategy is, of course, less compelling. Only the former truncation method guarantees that the EFT properly reproduces the behavior of UV models. We therefore suggest truncating the EFT series on amplitude level ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ for some $n\in\mathbb{N}$. As explicitly verified for Drell-Yan processes in Figs. 1 and 2 with $n=2$ and $4$ the ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-n})$ truncations (solid lines) provide better convergence toward the full model prediction than the cross-section truncation ${\mathcal{O}}_{\!\sigma}(\Lambda^{-n})$ (dashed lines). According to this prescription, the leading order ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-2})$ EFT contributions contain both interference of $d=6$ EFT operators with the SM and the $(d=6)^{2}$ contributions to the cross section, which can also be seen in Eq. (31). As discussed in Ref. Brivio:2022pyi , this provides a well-defined, unambiguous, and gauge-invariant setup. At the first subleading order ${\mathcal{O}}_{\!{\mathcal{A}}}(\Lambda^{-4})$ interference of $d=8$ operators with the SM, $(d=8)^{2}$ contributions, as well as the interference of $(d=6)$ with $(d=8)$ operators are included. In addition to dimension-eight operators contributing to the amplitude, the corresponding contributions from double insertion of two $d=6$ operators into the same amplitude are taken into account at this order as well, cf. Eq. (31), although this is not relevant for the present discussion of Drell-Yan tails at tree level.
@@ VI Illustration: $b\to c\tau\nu$
where we have again assumed down alignment for the left-handed quark doublet, and kept couplings to the $\tau$ lepton, and the second- and third-generation quarks. Other relevant constraints on this scenario are posed by the modifications of $Z$ and $W$ couplings to leptons induced by the RG evolution, which are sizable as they are proportional to the top Yukawa coupling $y_{t}$ in this case Feruglio:2016gvd ; Feruglio:2017rjo ; Cornella:2018tfd ; Allwicher:2022vbf ; Allwicher:2023aql . In our case, the biggest effect comes from $Z\nu\bar{\nu}$ vertex corrections, parametrized by Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga
@@
In summary, we find that EFTs can reliably describe Drell-Yan tails for $t$- and $u$-channel mediators with masses above $1.5$ TeV, including the models proposed to address the $R_{D^{(\ast)}}$ anomalies. We have shown that the EFT description at $\mathcal{O}_{\mathcal{A}}(\Lambda^{-2})$ is rather accurate for these masses and it can be systematically improved by including the $\mathcal{O}_{\mathcal{A}}(\Lambda^{-4})$ contributions.